OpenStudy (anonymous):

If a rock falls from a height of 33 meters on Earth, the height H (in meters) after x seconds is approximately H(x)=33-4.9x^2. When is the height of the rock 10 meters? (a) What is the height of the rock when x = 1.8 seconds? (b) When is the height of the rock 10 meters? (c) When does the rock strike the ground? I was able to get the answers for the first 2 questions. (a) is 17.124 and (b) 2.17. It's (c) that I'm still having trouble with.

OpenStudy (anonymous):

sorry that was not right set 4.9x^2=33 and solve for x

OpenStudy (anonymous):

Here's what I did: \[4.9x^2=33\] \[\frac{ 4.9x^2 }{4.9} = \frac{ 33 }{ 4.9 }\] \[\sqrt{x^2} = \sqrt{6.73}\] x = 2.59 I'm doing this homework on mymathlab and it says that my answer was wrong. I don't get what I did wrong.

OpenStudy (anonymous):

try 2.6

OpenStudy (anonymous):

because if you dont round the full answer comes out to 2.59513

OpenStudy (anonymous):

Oh, 2.6 worked. Thanks for the help!

OpenStudy (anonymous):

no problem